OpenStudy (anonymous):

] Suppose spin a spinner with eight equal sectors, numbered 1 to 8. What is the probability of spinning a 3 or an 8? _____________ 5. [1] True or False. Tom is late for class 50 times this semester. Therefore the probability he is late is 55%. This is an example of subjective probability. _____________ 6. [1] True or False. A discrete sample space is one in which outcomes are counted.________

OpenStudy (unklerhaukus):

\[P(3)=1/8\]\[P(8)=1/8\]\[P(3 \text{ or } 8)=P(3)+P(8)\]

OpenStudy (anonymous):

Which question is that for ?

OpenStudy (unklerhaukus):

4.

OpenStudy (anonymous):

so it would be 1/8+1/8?

OpenStudy (unklerhaukus):

yes and you can simplify a bit

OpenStudy (anonymous):

25%

OpenStudy (unklerhaukus):

good

OpenStudy (anonymous):

Thank you! so how would i do question 5?

OpenStudy (unklerhaukus):

i am not sure about 5. i dont think there is enough information given

OpenStudy (anonymous):

ok what about question 6? and there is one more question that i couldnt fit on the space provided

OpenStudy (anonymous):

true i guess

OpenStudy (anonymous):

i dont like these true false question can you help me with this question : The probability that a young Canadian plays soccer in the summer is 0.57. The probability that a young Canadian plays hockey in the winter is 0.83. Assuming that these are independent events, if a young Canadian is chosen at random, what is the probability that they play hockey or soccer? _____________

OpenStudy (unklerhaukus):

are those numbers percentages ?

OpenStudy (anonymous):

0.83 + 0.57 - 0.83* 0.57

OpenStudy (anonymous):

Yes

OpenStudy (unklerhaukus):

as in 83% or 0.83%

OpenStudy (anonymous):

83%